LATEX(1) LATEX(1) NAME latex,elatex,lambda,pdflatex−structuredtextformattingandtypesetting SYNOPSIS latex[firstline] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. TheLATEXlanguageisdescribedinthebookLATEX−ADocumentPreparationSystem.LATEXis aTEXmacropackage,notamodificationtotheTEXsourceprogram,soallthecapabilities describedintex(1)arepresent. TheLATEXmacrosencouragewriterstothinkaboutthecontentoftheirdocuments,ratherthan theform.Theideal,verydifficulttorealize,istohavenoformattingcommands(like‘‘switchto italic’’or‘‘skip2picas’’)inthedocumentatall;instead,everythingisdonebyspecificmarkup instructions:‘‘emphasize’’,‘‘startasection’’. TheprimarysourceofdocumentationforLATEXistheLATEXmanualreferencedbelow,andthe localguideinthefilelocal-guide.texorlocal.texorsomesuch. elatexisthee-TXextendedmodeversionofLATEXformat. lambdaistheOmegaversionoftheLATEXformat. pdflatexisthepdfTEXversionoftheLATEXformat. Onsomesystemslatex209andslitexareavailableforcompatibilitywitholderversionsof LATEX.Theseshouldnotbeusedfornewtexts. SEEALSO amslatex(1),amstex(1),tex(1), LeslieLamport,LATEX−ADocumentPreparationSystem,Addison-Wesley,1985,ISBN 0-201-15790-X. Web2C7.32x 29March1999 1